Last summer, Kyle opened Subterranean: A New Kind of Sandwich Shop in Akron, Ohio. The following is his production schedule Subterranean: A New Kind of Sandwich Shop Subs Total Workers Ovens per Oven week Cost Labor Food TC ATC MPL ATC AFC AVC MC Cost Cost 2 0 1000 0 0 N/A N/A N/A N/A N/A 2 150 1000 200 300 2 2 400 1000 400 800 2 600 1000 600 1200 2 700 1000 800 1400 2 750 1000 1000 1500 2 770 1000 1200 1540 2 775 1000 1400 1550 What is the average variable cost of the fourth row?
Answer #1

4th row means output from 3 workers.

Sandwich produced = 600

FC = 1000

TVC = 600 + 1200 = 1800

AVC = TVC/Q = 1800/600 = 3